OpenStudy (anonymous):

Simplify without negative exponents. y^-5 / y^-2

OpenStudy (uri):

without negative exponents..? :o

OpenStudy (anonymous):

yes im unsure about how to do this problem

OpenStudy (unklerhaukus):

\[\frac{x^m}{x^n}=x^{m-n}\]

OpenStudy (anonymous):

y^5 / y^2 = y^5-2= y^3 ???

OpenStudy (unklerhaukus):

yes. but that is not that same as the original question

OpenStudy (anonymous):

it says simplify without negatives?? y^-5 / y^-2 = y^-5--2= y^-3

OpenStudy (unklerhaukus):

thats right

OpenStudy (unklerhaukus):

you can change it a bit \[x^{-n}=\frac1{x^n}\]

OpenStudy (anonymous):

okay thanks !